IN TRIANGLE ABC CotA : Cot B : Cot C = 30;16;9, FIND THE RATIO a;b;c = what is the ratio

Grade:

2 Answers

sushant subhash rajbhar
11 years ago

dear frnd!!!

u first get it into sin and cos !! and after den u can apply the sine formula   :) :)

Kaushik Kumar
11 years ago

In a Triangle 

Cot A*Cot B + Cot B*CotC + Cot C*Cot A = 1

Now, Cot A/30=Cot B/16=Cot C/9 = x (say)

put it above

x^2*(30*16+16*9+9*30) = 1

x^2 = 1/894

x     = y(say)

therefore , Cot A = 5/y , Cot B = 16/y , Cot C = 9/y

now get ,

sin A = (1+cot A^2)^1/2

sin B = (1+cot B^2)^1/2

sin C = (1+cot C^2)^1/2

required ratio .